Problem 6: For the foundation shown below with eL = 1.2 ft, eB = 0.15 ft, calculate the allowable load that the foundation can carry using a factor of Safety of 3 and the General Bearing Capacity.Problem б: For the foundation shown below with e-= 1.2 ft. ев = 0.15 ft, calculate the allowable load that the foundation can carry using a factor of Safety of 3 and the General Bearing Capacity oblein Sand d=35° c=0 γ = 115 lb/ft3 4 ft x 6 ft
Answer #1

в- В- аеь L= L- 2e. %3D 2(0.15) = 6 - 2(1-2) - 3. 6 ft 8 = 3.1 ft 13.32 ft Qsafe = Ax Vs %3D (1+0.3Je+ y-1) + •.53 (1-02) V